OpenStudy (anonymous):

Calculate the discriminant. x2 – 2x + 4 = 0 A. –12 B. 20 C. 12 D. 0

OpenStudy (michele_laino):

the requested discriminant is given by the subsequent computation: \[\Delta {\text{ = }}{\left( { - 2} \right)^2} - 4 \times 1 \times 4 = ...?\]

OpenStudy (anonymous):

okay!